TheJag18:

(Comparing Types of Interest MC) A college student is paying for tuition through private loans. Two lenders have approved the student for a $25,000 loan. Offer 1: 5.99% annual simple interest, with a total account balance of $32,487.50 after a 60-month term Offer 2: 3.75% annual interest compounded monthly for a 66-month term Assuming no payments are made, what is the difference in the account balances at the end of the loan terms? Round your answer to the nearest penny. $1,245.00 $1,770.87 $2,964.36 $3,319.94

TheJag18:

you have the work?

TheJag18:

@sza

toga:

we can use the simple interest formula to calculate the balance for Offer 1: I = P * r * t I = 25,000 * 0.0599 * 5 I = 7,475 Total account balance for Offer 1 = P + I Total account balance for Offer 1 = 25,000 + 7,475 Total account balance for Offer 1 = 32,475 For Offer 2, we can use the compound interest formula: A = P * (1 + r/n)^(n*t) A = 25,000 * (1 + 0.0375/12)^(12*66/12) A = 31,692.17 The difference in account balances at the end of the loan terms is: Difference = Total account balance for Offer 2 - Total account balance for Offer 1 Difference = 31,692.17 - 32,487.50 Difference = -795.33 Rounding to the nearest penny, the difference in account balances at the end of the loan terms is $-795.33. This means that Offer 1 would result in a higher account balance at the end of the loan term.